# mO shizhe[#1] VeraCT Scan: Retrieval-Augmented Fake News Detection with Justifiable Reasoning

### v1

@@ 1 Introduction
Traditional approaches in fake news detection have primarily focused on the linguistic features, also called content-based detection (Castillo et al., 2011; Pérez-Rosas et al., 2017; Giachanou et al., 2019; Przybyla, 2020; Giachanou et al., 2020; Sheikhi, 2021; Kirchknopf et al., 2021; Zhou et al., 2020), which demands laborious feature engineering and is ineffective when the fake news is written by imitating the real news to mislead intentionally. Another line of research is the social context-based method (Qazvinian et al., 2011; Baly et al., 2018; Shu et al., 2019b; Monti et al., 2019; Nan et al., 2023), which analyzes the interactions among users, publishers, and posts. However, the feasibility of obtaining user information is challenging for the real-world application. A more recent research approach is the knowledge-based method (Hu et al., 2021; Saeed et al., 2022; Pan et al., 2023; Chen et al., 2023b; Liao et al., 2023; Zhang and Gao, 2023; Li et al., 2023), which discerns the veracity of the claim by comparing against the evidence retrieved from external knowledge base. However, current approaches often do not fully utilize external resources like the Internet. Additionally, there is a lack of development and optimization of a comprehensive end-to-end pipeline that includes news comprehension, search optimization, verification, and reasoning.
In this paper, we introduce VeraCT Scan, a novel retrieval-augmented system for fake news detection. VeraCT Scan initiates this process by delineating key facts across multiple levels of granularity. For each identified fact, a comprehensive internet search is conducted to gather relevant information. Then, the veracity of the news is determined by combining this typically disparate and conflicting information, taking into account the varying degrees of source credibility. To increase the trustworthiness of our approach, we underscore the necessity of a transparent reasoning process and provide rationales for each supporting or conflicting judgment.
@@ 2.1 Fake News Detection
Existing fake news detection methods can be categorized into three types: 1) Content-Based Methods (Sheikhi, 2021; Pérez-Rosas et al., 2017; Castillo et al., 2011; Przybyla, 2020; Giachanou et al., 2019; Huang et al., 2023; Giachanou et al., 2020; Kirchknopf et al., 2021; Nakamura et al., 2020; Chen et al., 2023b; Zhou et al., 2020) which analyze articles’ linguistic features (e.g., text length, punctuation usage, emotion symbols) to differentiate fake news from real ones. However, these methods demand laborious feature engineering and are often ineffective when fake news is written to intentionally mislead readers. 2) Social Context-Based Methods (Shu et al., 2019b; Nan et al., 2023; Baly et al., 2018; Monti et al., 2019; Qazvinian et al., 2011) which analyze the interactions among users, publishers, and posts to detect fake news. However, the feasibility of obtaining user information in the news propagation process presents challenges for the real-world applicability of this method. 3) Fact-Based Methods (Saeed et al., 2022; Pan et al., 2023; Hu et al., 2021; Xu et al., 2023; Chen et al., 2023b; Cheung and Lam, 2023) which focus on claim verification by comparing against external knowledge. These methods fall short in providing an end-to-end solution that considers information seeking and the management of conflicting evidence.
Recently, Wang and Shu (2023) leverage large language models (LLMs) to decompose complex claims into sequences of first-order logic, and then guide the search and information verification. Different from their work, to classify fake news, we propose a pipeline that is expanded to incorporate additional steps for extracting facts from articles. Liao et al. (2023) outlines a multi-step process for detecting fake news, which consists of news summarization, searching, and verification. In contrast to their method, we employ LLMs instead of specifically trained encoder-decoder transformers for these natural language processing tasks. In addition, we leverage source credibility to differentiate conflicting evidences, a common challenge in real-world news verification that has rarely been explored in previous research.
@@ 2.2 Retrieval-Augmented Generation
The integration of retrieval-augmented generation (RAG) allows LLMs to extend beyond the limits of the training corpus by retrieving information from external knowledge bases before the generative process (Lewis et al., 2020; Chen et al., 2023a). RAG has emerged as a solution to overcome the limitations of LLMs including the challenge of out-of-date knowledge and the tendency to produce hallucinations or irrelevant and factually incorrect content. By integrating external, up-to-date documents into the generation process, LLMs can generate more reliable responses across a broad spectrum of tasks, including open-domain question answering (Izacard and Grave, 2020; Trivedi et al., 2022; Li et al., 2022; Xu et al., 2024), dialogue systems (Cai et al., 2018; Peng et al., 2023), and code generation (Zhou et al., 2022). RAG is also commonly integrated into commercial chatbot products to provide updated information, e.g Perplexity222https://www.perplexity.com and Gemini333https://gemini.google.com. In this paper, we leverage RAG for fake news detection by generating both verdicts and justifications.
@@ 3 Approach
Figure 1 shows the main workflow of VeraCT Scan. We prompt GPT-4 Turbo for key fact extraction, query generation, verification, and rationale generation (See Appendix A for prompts being used). The outputs from GPT-4 Turbo, supplemented with manual reviews, serve as training data to fine-tune Llama-2 13B Touvron et al. (2023), enabling it to support these tasks as well. Regarding the search component, we employ both Google and our proprietary in-house news search engine for comprehensive information retrieval.
@@ 3.4 Source Credibility and Final Decision
When researching a given topic, it is common to encounter conflicting information on the Internet. Therefore, assessing the credibility of each information source is crucial. mediabiasfactcheck.com is one of the most comprehensive resources for media bias on the Internet, offering credibility ratings for over 7,600+ news publishers. Besides, NewsBreak also developed a proprietary 5-level credibility ratings for 30,000+ publishers.
Both rating systems serve as features to train a LightGBMKe et al. (2017) classifier that determines the likelihood of a fact claim being true. Besides, domain and verification flags (i.e. support, negate, or baseless) from each search result are also used as classification features.
@@ 3.5 Llama-2 13B Fine Tuning
Following previous studiesZhou et al. (2023); Taori et al. (2023), we utilize a mixed dataset of diverse tasks for supervised fine-tuning (SFT). Outputs of GPT-4 Turbo from the tasks described above are used as part of the training data. Specifically, we purposely modify some key claims being extracted from news articles into fake ones when generating claim verification data set. Besides, the following datasets have also been incorporated into the training set:
@@
QA with RAG: GPT-4 generated answers to questions in NewsBreak search using knowledge retrieved from our proprietary search engine.
@@
WebGLMLiu et al. (2023a): web-enhanced question-answering dataset.
@@
To enhance the capability of processing long inputs, we trained the model with RoPE scalingSu et al. (2021); Liu et al. (2023b). Specifically, we adjusted the context window size in SFT to be twice as large as that in the original Llama-2 model, setting it to 8192 tokens, and we set the scaling factor at 2.0. We employed full training with an initial learning rate of 1e-5, and limited the training to 1 epoch. The training process was executed on four NVIDIA A100 GPUs.
@@ 4.1 Datasets
This dataset consists of news articles shared on Facebook during the week surrounding the 2016 U.S. election. It includes data collected from nine different news agencies, spanning from September 19 to 23, and then September 26 and 27. Each article was fact-checked by a team of five BuzzFeed journalists. The articles are categorized under four labels: mostly true, mostly false, a mix of true and false, and no factual content. In line with Shu et al. (2019a), we utilize the subset of 182 news articles for our benchmark. Each article in this subset has been assigned one of two binary labels (true or fake news), making it suitable for our binary classification setting.
@@
This dataset comprises a total of about 1,200 claims from Politifact and Snopes777https://www.snopes.com, including both true and false claims. It distinguishes itself from other publicly available datasets by featuring only recent claims spanning from 2022 to March 2024. We will release this dataset upon the publication of this paper.
@@ 4.3 Implementation Details
To aid in the verification of news articles, the main claim of each news article is identified and then compared against internet search results. To ensure a fair comparison, we have developed heuristics to carefully filter out fact-checking content from search engine results in all the experiments below.
The datasets above except LLMFake are each aggregated to train the final classifier, utilizing the features outlined in Section 3.4, and subsequently report the end-to-end accuracy. Both the training and testing processes are conducted using a 5-fold cross-validation approach. We also provided baseline benchmarks for comparison.
@@ 5 Experimental Results
| GPT-4-based Zero-shot Detector (COT) (Chen and Shu, 2023) | ||||

### v2

@@ 1 Introduction
Traditional approaches in fake news detection have primarily focused on the linguistic features, also called content-based detection (Castillo et al., 2011; Pérez-Rosas et al., 2018; Giachanou et al., 2019; Przybyla, 2020; Giachanou et al., 2020; Sheikhi, 2021; Kirchknopf et al., 2021; Zhou et al., 2020), which demands laborious feature engineering and is ineffective when the fake news is written by imitating the real news to mislead intentionally. Another line of research is the social context-based method (Qazvinian et al., 2011; Baly et al., 2018; Shu et al., 2020; Monti et al., 2019; Nan et al., 2023), which analyzes the interactions among users, publishers, and posts. However, the feasibility of obtaining user information is challenging for the real-world application. A more recent research approach is the knowledge-based method (Hu et al., 2021; Saeed et al., 2022; Pan et al., 2023; Chen et al., 2023; Liao et al., 2023; Zhang and Gao, 2023; Li et al., 2024), which discerns the veracity of a factual claim by comparing against the evidence retrieved from external knowledge base. However, current approaches often do not fully utilize external resources like the Internet. Additionally, there is a lack of development and optimization of a comprehensive end-to-end pipeline that includes news comprehension, search optimization, verification, and reasoning.
In this paper, we introduce VeraCT Scan, a novel retrieval-augmented system for fake news detection. VeraCT Scan initiates this process by identifying key factual claims across multiple levels of granularity. For each identified factual claim, a comprehensive internet search is conducted to gather relevant information. Then, the veracity of the news is determined by combining this typically disparate and conflicting information, taking into account the varying degrees of source credibility. To increase the trustworthiness of our approach, we underscore the necessity of a transparent reasoning process and provide rationales for each supporting or conflicting judgment.
@@ 2.1 Fake News Detection
Existing fake news detection methods can be categorized into three types: 1) Content-Based Methods (Sheikhi, 2021; Pérez-Rosas et al., 2018; Castillo et al., 2011; Przybyla, 2020; Giachanou et al., 2019; Huang et al., 2023; Giachanou et al., 2020; Kirchknopf et al., 2021; Nakamura et al., 2020; Chen et al., 2023; Zhou et al., 2020) which analyze articles’ linguistic features (e.g., text length, punctuation usage, emotion symbols) to differentiate fake news from real ones. However, these methods demand laborious feature engineering and are often ineffective when fake news is written to intentionally mislead readers. 2) Social Context-Based Methods (Shu et al., 2020; Nan et al., 2023; Baly et al., 2018; Monti et al., 2019; Qazvinian et al., 2011) which analyze the interactions among users, publishers, and posts to detect fake news. However, the feasibility of obtaining user information in the news propagation process presents challenges for the real-world applicability of this method. 3) Fact-Based Methods (Saeed et al., 2022; Pan et al., 2023; Hu et al., 2021; Xu et al., 2023; Chen et al., 2023; Cheung and Lam, 2023) which focus on factual claim verification by comparing against external knowledge. These methods fall short in providing an end-to-end solution that considers information seeking and the management of conflicting evidence.
Recently, Wang and Shu (2023) leverage large language models (LLMs) to decompose complex claims into sequences of first-order logic, and then guide the search and information verification. Different from their work, we propose a pipeline that includes full steps to classify fake news. Liao et al. (2023) outlines a multi-step process for detecting fake news, which consists of news summarization, searching, and verification. In contrast to their method, we employ LLMs instead of specifically trained encoder-decoder transformers for these natural language processing tasks. In addition, we leverage source credibility to differentiate conflicting evidences, a common challenge in real-world news verification that has rarely been explored in previous research.
@@ 2.2 Retrieval-Augmented Generation
The integration of retrieval-augmented generation (RAG) allows LLMs to extend beyond the limits of the training corpus by retrieving information from external knowledge bases before the generative process (Lewis et al., 2020; Chen et al., 2024). RAG has emerged as a solution to overcome the limitations of LLMs including the challenge of out-of-date knowledge and the tendency to produce hallucinations or irrelevant and factually incorrect content. By integrating external, up-to-date documents into the generation process, LLMs can generate more reliable responses across a broad spectrum of tasks, including open-domain question answering (Izacard and Grave, 2021; Trivedi et al., 2023; Li et al., 2023; Xu et al., 2024), dialogue systems (Cai et al., 2019; Peng et al., 2023), and code generation (Zhou et al., 2023b). RAG is also commonly integrated into commercial chatbot products to provide updated information, e.g Perplexity222https://www.perplexity.com and Gemini333https://gemini.google.com. In this paper, we leverage RAG for fake news detection by generating both verdicts and justifications.
@@ 3 Approach
In this paper, the term "claim" refers to the fact stated in a news article. The terms "factual claim extraction" and "fact extraction" are used interchangeably throughout the paper.
Figure 1 shows the main workflow of VeraCT Scan. We prompt GPT-4 Turbo for key fact extraction, query generation, verification, and rationale generation (See Appendix A for prompts being used). These individual components can be easily exchange to other LLMs or search engines. In this work, the outputs from GPT-4 Turbo, supplemented with manual reviews, serve as training data to fine-tune Llama-2 13B Touvron et al. (2023), enabling it to support these tasks as well. Regarding the search component, we employ both Google and our proprietary in-house news search engine for comprehensive information retrieval.
@@ 3.4 Source Credibility and Final Decision
When researching a given topic, it is common to encounter conflicting information on the Internet. To avoid bias from single source, multiple sources are used to corroborate each other. Therefore, assessing the credibility of each information source is crucial. Mediabiasfactcheck.com is one of the most comprehensive resources for assessing media bias on the internet, offering credibility ratings for over 8,000 news publishers. Similarly, NewsBreak has developed a proprietary 5-level credibility rating system for more than 30,000 publishers. While NewsBreak’s ratings are also based on the credibility of source content, unlike mediabiasfactcheck.com, NewsBreak does not identify the political bias of the sources.
In this paper, NewsBreak’s rating systems serves as features to train a LightGBMKe et al. (2017) classifier that determines the likelihood of a fact claim being true. Besides, domain and verification flags (i.e. support, negate, or baseless) from each search result are also used as classification features.
@@ 3.5 Llama-2 13B Fine Tuning
Following previous studiesZhou et al. (2023a); Taori et al. (2023), we utilize a mixed dataset of diverse tasks for supervised fine-tuning (SFT). Outputs of GPT-4 Turbo from the tasks described above are used as part of the training data. Specifically, we purposely modify some key factual claims being extracted from news articles into fake ones when generating claim verification data set. Besides, the following datasets have also been incorporated into the training set:
@@
QA with RAG: GPT-4 generated answers to questions in NewsBreak search logs using knowledge retrieved from our proprietary search engine.
@@
WebGLMLiu et al. (2023): web-enhanced question-answering dataset.
@@
To enhance the capability of processing long inputs, we trained the model with RoPE scalingSu et al. (2021); Liu et al. (2024). Specifically, we adjusted the context window size in SFT to be twice as large as that in the original Llama-2 model, setting it to 8192 tokens, and we set the scaling factor at 2.0. We employed full training with an initial learning rate of 1e-5, and limited the training to 1 epoch. The training process was executed on four NVIDIA A100 GPUs.
@@ 4.1 Datasets
This dataset consists of news articles shared on Facebook during the week surrounding the 2016 U.S. election. It includes data collected from nine different news agencies, spanning from September 19 to 23, and then September 26 and 27. Each article was fact-checked by a team of five BuzzFeed journalists. The articles are categorized under four labels: mostly true, mostly false, a mix of true and false, and no factual content. In line with Shu et al. (2019), we utilize the subset of 182 news articles for our benchmark. Each article in this subset has been assigned one of two binary labels (true or fake news), making it suitable for our binary classification setting.
@@
The dataset was manually collected from the prestigious fact-checking organizations PolitiFact and Snopes777https://www.snopes.com. It includes approximately 1,200 verifiable claims along with the fact-check rating labels that determine the level of truthfulness for each claim. The clarifications for the labels and the additional detailed analysis reports were not collected. Non-text-based claims were filtered out, and exclusive fact-checks with supporting sources specific to these organizations were also filtered out.
@@ 4.3 Implementation Details
To aid in the verification of news articles, the main factual claim of each news article is identified and then compared against internet search results. To ensure a fair comparison, we have developed heuristics to carefully filter out fact-checking content from search engine results in all the experiments below.
The datasets above except LLMFake are each aggregated to train the final LightGBM classifier, utilizing the features outlined in Section 3.4, and subsequently report the end-to-end accuracy. Both the training and testing processes are conducted using a 5-fold cross-validation approach. We also provided baseline benchmarks for comparison.
@@ 5 Experimental Results
| GPT-4-based Zero-shot Detector (COT) (Chen and Shu, 2024) | ||||
@@ 6 Conclusion and Future Work
Furthermore, we observe that within the entire system, the majority of errors occur during the verification stage, with a smaller fraction arising during the claim extraction phase. The causes of these errors include: (i) Irrelevant search results used for verification. (ii) Updated news events leading to outdated reports being used for verification. (iii) Each report only supporting a part of the claim, necessitating the proper merging of relevant information from multiple news reports for full verification. (iv) Improper normalization of named entities or temporal expressions during the claim extraction stage, making alignment difficult during verification (e.g., "last weekend" vs. an exact date). We hope to address these issues in future work.